MODS1 AGw Commissioning at LDG: 19 November 2010 UTC

Observer(s): R. Pogge and B. Shappee (OSU), M. Pedani and M. Wagner (LBTO)
Telescope Operator: G. Bechetti (LBTO) with D. Gonzalez in Tucson
Active Optics Support: J. Hill (LBTO) and D. Miller (LBTO remote from Tucson)
Telescope Support: J. Little
Instrument Support: D. Officer (LBTO)


0055 - We Open. Weather: T_air = 4.6 RH ~ 5% Wind ~6 m/s Clear Sky T_mirror = 3.0

Seeing looks to be beautiful - in binoculars Kitt Peak was crystal clear and rock steady when backlit by the sunset. We could see the 4m, 90-inch, and the McMath triangle

Best IQ ever seen with MODS to date.

A few issues: Andrew Rakich found a problem with the WFS analysis code that will be fixed tomorrow. It leads to growing aberrations, especially astigmatism and coma, particularly if the WFS/guide star is relatively far off-axis. We tried his recommendation of choosing less off-axis guide stars and it worked well, but see notes about some of the manifestations of this problem before he suggested a way to mitigate it. They found the problem and should have it implemented in the GCS for tomorrow.

Ran with the first-cut AGw-to-SFP transformations from last night. Some refinement is needed and we got a great run on the Stone E field in good (0.5-0.7 arcsec) seeing to accomplish this. Will evaluate the data tomorrow.

Ran a full set of 1- and 10-arcsec 5-Dice sequences at elevations of 30 to 78 degrees. Will analyze tomorrow.

Given the excellent seeing, we took MODS for a little spin to see what it could do under such circumstances. It was a very nice demonstration that if LBT throws MODS excellent (0.35-0.4 arcsec) seeing, it can use it. Image quality in red was excellent, spectral quality in blue was good even down at 370nm were with a 0.3-arcsec slit and the grating we successfully and cleanly resolved the [O II]3727,29A doublet line in a planetary nebula.

A great night!


2330 - MODS dewars filled - hold time still excellent.


0105 - Pointing Check

    IE = 11.0
    CA = 12.0

0115 - Collimation Check on BS918. Seeing ~ 0.7-0.8" at El.= 56

0140 Seeing ~ 0.5" on MODS smile

0141 Slew to Stone_O to start Collect Transformations data for the first 11 stars at PA=0.0

Stone_O Field:
1   0
11  0
15  0
24  0
23  0
28  0
26  0
33  0
32  0
12  0
10  0

We test the AGW Transformations as derived from yesterday's Stone-E

The wfs images are in /OldRepository/AGW_Data/20101119/

Starting with wfsc000001 and mods1r.20101119.0002

As yesterday Star 24 probably acquired a neighboring star.

0201 - Seeing on MODS images is 0.45" (new record for MODS with Rick jumping up on the chair). Image: mods1r.20101119.0004

0209 - Last Images: left_wfsc000015.fits and mods1r.20101119.0012.fits

Good Seeing Images and Spectra

Getting some opportunistic good-seeing images and spectra before we have to dive into the main list of tasks for the night - you don't get conditions like this often, so let's see what MODS can do...

0210 - Slewing to M27 for nice "ugriz' pictures + some spectra

0300 - Looking at a spectrophotometric standard in the 1.2-arcsec slit - in this great seeing, we can get a good throughput calculation before our fat spectrophotometer slit is made, and we get to exercise the acquisition sequence.

Target: Feige 110 - 10, 60 acquisitions and 3x300s exposure in both channels. Tons of signal in both channels.

Images in the acquisition snapshot are funny looking with a long tail. See particularly image mods1r.20101119.0025, and in the subsequent blue and red spectra. The "tail" is aligned fortuitously with the slit, and its appearance in both channels indicates it is in the telescope and not the instrument. It was not there in previous images, and Marco later noticed there was a change in some of the primary mirror collimation values between 0024 and 0025. We've made a note to Andrew and Doug about this.

Then shot Ne, Ar, and Hg lamp spectra for wavelength calibration after stopping guiding. This gave Geno a chance to reset the mirrors which were getting close to limits. Mark W. will analyze these to make a first cut at a spectroscopic throughput estimate in grating mode.

Five Dice Offset Tests

0429 - Slewing to Stone-A2 for Five-Dice measurements at El=55-degrees

0443 - Script failed because of the preset spontaneously canceled and we lost the guide/wfs view between offset 4 and offset 5 of the first 5-dice pattern.

We resent the preset and started over again at 0445. David is watching to see what happened. Mark Wagner noticed at about the same time as we saw the guide screens blank out that the network appeared to hang briefly. Not clear if it is related.

0513 - We accidentally stalled a script and restarted. Moonlight is an issue in these fields, and seeing is now a less-spectacular 0.6-0.7, but we note that the humidity has increased from 10% to 40% and plateaued in the last hour, and the wind is now up around 5m/s.

Ran 10- and 1-arcsec patterns at this elevation.

0610 - Field at El=75 degrees, 10- and 1-arcsec patterns

0710 - Field at El=30 degree, 10-arcsec pattern

0750 - Getting too low for a 1-arcsec, looking for another El=30 field that is rising instead of setting.

Wind is coming up to 10m/s sustained, so closing vent doors. Seeing will not be getting much better as a consequence.

0809 - Starting 1-arcsec 5-dice at El=30 and rising

0851 - Starting 10-arcsec 5-dice at El=63 and rising

0914 - Weather update: T_air = +1 .5, RH=46%, Wind ~ 10m/s, T_mirror = +1.4, Seeing~0.8" in the AGW

0920 - Starting 1-arcsec 5-dice at El = 57 and rising

0928 - We see the stars developing a "tail".

Images: mods1r.20101119.0250 (UT: 09:24:09.282) to 0254 (UT: 09:28:02.898)

We select another star more on-axis and ...

Image: mods1r.20101119.0255 (UT: 09:32:46.176) is very good with 0.53" FWHM.

UT Time Elev. M1 X M1 Y M1 Z M1 ROTX M1 ROTY Image name
09:24:09.282 57.26879 -0.863 -0.026 0.674 -8.860 -5.495 /Repository/20101119/mods1r.20101119.0250.fits
09:25:07.343 57.32407 -0.869 -0.050 0.673 -9.121 -5.444 /Repository/20101119/mods1r.20101119.0251.fits
09:26:05.185 57.37777 -0.860 -0.053 0.676 -9.144 -5.563 /Repository/20101119/mods1r.20101119.0252.fits
09:27:04.161 57.43007 -0.894 -0.047 0.677 -9.076 -5.231 /Repository/20101119/mods1r.20101119.0253.fits
09:28:02.898 57.48118 -0.891 -0.030 0.678 -8.912 -5.265 /Repository/20101119/mods1r.20101119.0254.fits
09:32:46.176 57.70599 -1.047 -0.680 0.609 -15.879 -3.673 /Repository/20101119/mods1r.20101119.0255.fits

0957 We are done with 5-Dice tests

0958 - Check Pointing parameters

    IE = 5.0
    CA = 24.0

Good Seeing Multi-Color Picture

Looks good enough to go for a nice "first light" shot in imaging while the folks downtown are preparing for the next transformation field...

1010 - M1 , ugriz sequence using photcal. Seeing is 0.5-arcsec or better

Stone E Transformation field.

1023 - Slew to StoneE for first set of Transformations Data (18 stars).

transform_collect, 'Stone_E.list', /MODS, START=1


41   -90
45   -90
106  -90
55   -90
68   -90
24   -90
1   -90
16   -90
29   -90
40   -90
57   -90
119   -90
86   -90
63   -90
121   -90
81   -90
67   -90
85   -90

First WFS image: /OldRepository/AGW_Data/20101119/left_wfsc000016.fits
First MODS Image: mods1r.20101119.0288.fits

Last WFS image: /OldRepository/AGW_Data/20101119/left_wfsc000036.fits
Last MODS Image: mods1r.20101119.0305.fits

1103 - transform_collect, 'Stone_Ep00.list', /MODS, START=1


30  0
5   0
9   0
24  0
34  0
31  0
60  0
51  0
54  0
71  0
72  0
112 0
102 0
128 0
136 0

Star #102 (left_wfsc000050.fits) has a companion!

First WFS image: /OldRepository/AGW_Data/20101119/left_wfsc000037.fits
First MODS Image: mods1r.20101119.0306.fits

Last WFS image: /OldRepository/AGW_Data/20101119/left_wfsc000053.fits
Last MODS Image: mods1r.20101119.0320.fits

11:33 - transform_collect, 'Stone_Ep90.list', /MODS, START=1


31  90
30  90
50  90
66  90
93  90
124 90
116 90
117 90
20  90

First WFS image: /OldRepository/AGW_Data/20101119/left_wfsc000054.fits
First MODS Image: mods1r.20101119.0321.fits

Last WFS image: /OldRepository/AGW_Data/20101119/left_wfsc000065.fits
Last MODS Image: mods1r.20101119.0329.fits

11:54 - Transformations done

Another spectral target, and some test calibrations

12:00 - Acquires Akn120 at elevation 33-degrees. No problems acquiring, took 3 minute (by hand) from image to start of spectrum.

The offsetting is accurate and nails the object on the slit. Guiding keeps it locked solidly on the slit for 2x300s exposures, even at El=30 where the guide star is getting elongated. WFS kept up pretty well.

1230 - While Geno is closing up, trying a few spectral lamps and flats, but not too many as we're beat, and it gives the instrument something to do while we're closing out the night report.

-- MarcoPedani - 19 Nov 2010
Topic revision: r11 - 16 Aug 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback